OpenStudy (anonymous):

using definition of a null sequence prove that a_n=(1+(-1)^n)/n+(-1)^n, n=2,3,... is null

OpenStudy (anonymous):

\[\lim _{n \rightarrow \infty} a _{n}=\lim _{n \rightarrow \infty} (1+(-1)^{n})/n+(-1)^{n}\] clearly the numerator is 2,0,2,0,2,0,2,0....... and the denominator is growing infinitly. So the a_n is getting smaller and smaller. By definition \[|a _{n}-0| =|a _{n}|<\epsilon \] since lim a_n is 0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

use what i wrote befor

OpenStudy (anonymous):

by definition, you have to find n big enough so that for this n \[|a _{n}-0|=|a _{n}|<\epsilon\]

OpenStudy (anonymous):

\[|1+(-1)^{n}/n+(-1)^{n}|<\]

OpenStudy (anonymous):

epsilon missing

OpenStudy (anonymous):

for odd n |a_n| is 0. so just do it for even n: 2/n+1<e 2/e-1<n

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you wellcome

OpenStudy (anonymous):

ah sorry new to all this, thanks